Consider the following abbreviated financial statements for Rarrothead Enterprises: PARROTHEAD ENTERPRISES PARROTHEAD ENTERPRISES 2014 and 2015 Partial Balance Sheets 2015 Income Statement Liabilities &OwnersSales Assets 12590 2015 2014 2015 2014, Costs 5594 Current assets 03900abilities 325 408 Depreciation 1035 Net fixed assets 3642 4138 Long-term debt 2019 2135 Interest 166 What is the change in net working capital for 2015?
Answer #1

CHANGE IN NET WORKING CAPITAL FOR 2015:

1. NET WORKING CAPITAL= CURRENT ASSETS - CURRENT LIABILITIES

2. NET WORKING CAPITAL FOR 2014 = (900-408)=492

3.NET WORKING CAPITAL FOR 2015 = (903-325)=578

4. CHANGE IN NET WORKING CAPITAL=(3-2) = (578-492)= 86